The output of a process is valued at $100 per unit. The cost of labor is $50 per hour including benefits. The accounting department provided the following information about the process for the past four weeks:

Week 1

Week 2

Week 3

Week 4

Units produced

1,124

1,310

1,092

981

Labor ($)

12,735

14,842

10,603

9,526

Material ($)

21,041

24,523

20,442

18,364

Overhead ($)

8,992

10,480

8,736

7,848

a.   Use the multifactor productivity ratio to see whether recent process improvements had any effect and, if so, when the effect was noticeable.

b.   Has labor productivity changed? Use the labor productivity ratio to support your answer.
Answer #1

Total value of unite produced:-           

For week 1=1124*$100=$112400

For week 2=1310*$100=$131000

For week 3=1092*$100=$109200

For week 4=981*$100=$98100

Total Labour hours:-

For week 1=12735/$50=254.7

For week 2=14842/$50=296.8

For week 3=10603/$50=212.1

For week 4=9525/$50=190.5

Answer A:-

Multifactor productivity ratio

For week 1= output/input =112400/42768=2.628

For week 2= output/input =131000/49845=2.628

For week 3= output/input =109200/39781=2.745

For week 4= output/input =98100/35738=2.745

Difference = (2.745-2.628/2.628)*100 =4.45%

Thus there is an improvement in Multifactor productivity ratio in week 3.

Answer B:-

Labour productivity:-

For week 1=output/input =1124/254.7 =4.413/hour

For week 2=output/input =1310/296.84 =4.413/hour

For week 3=output/input =1092/212.06 =5.1495/hour

For week 4=output/input =981/190.52 =5.1491/hour

Changes =(5.1491-4.4130/4.413)*100% =16.68%

So labour productivity improved by 16.68%

KINDLY RATE THE ANSWER AS THUMBS UP. THANKS A LOT.